OpenStudy (anonymous):

Given the function g(x) = -4x + 5, find g(-4). Please. and thank you.

OpenStudy (anonymous):

@jim_thompson5910 Please.

jimthompson5910 (jim_thompson5910):

replace x with -4 and evaluate

jimthompson5910 (jim_thompson5910):

g(x) = -4x + 5 g(-4) = -4(-4) + 5

jimthompson5910 (jim_thompson5910):

what do you get when you do?

OpenStudy (anonymous):

-4g = 16 + 5?

jimthompson5910 (jim_thompson5910):

oh, i guess it's confusing notation...sry I should have said that g(-4) means "the function value at -4"...it's not the same as g times -4

jimthompson5910 (jim_thompson5910):

so it should be g(-4) = 16+5

OpenStudy (anonymous):

so what is the answer?

jimthompson5910 (jim_thompson5910):

just add 16 and 5 to get it

OpenStudy (anonymous):

oh.

jimthompson5910 (jim_thompson5910):

so g(-4) = 21

OpenStudy (anonymous):

so the answer is 21

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

kk thanks.